# Energy landscape interpretation of universal linearly increasing absorption with frequency

### v2

@@ Abstract
Absorption of elastic waves in complex media often depends on frequency in a linear way for both longitudinal and shear waves. This universal property occurs in media such as rocks, unconsolidated sediments, and human tissue. Absorption is due to relaxation processes at atomic scale up to the sub-micron scale of biological materials, and we argue that these processes are thermally activated. Unusual for ultrasonics and seismics, we can therefore express absorption as an integral over an activation energy landscape weighted by an energy distribution. The universal power-law property corresponds to a flat activation energy landscape corresponding to maximal randomness.
@@ I Introduction
The amplitude of elastic waves undergoes attenuation which often follows a power-law in frequency, $\omega$:
@@
where ${u}$ is particle velocity and $x$ is distance. The two mechanisms for attenuation, $\alpha$, are viscoelastic absorption leading to heating, and scattering of energy. This paper is only concerned with mechanisms for absorption, although medium properties may also be inferred from the scattering, [1]. In complex media absorption often increases linearly with frequency, i.e. $y=1$, [2, Sect. 5.1], as the many examples that follow demonstrate.
In seismology, many nuclear-explosion and earthquake data sets in the range $10^{-3}$ to 10 Hz appear to have a constant quality factor, $Q$, for shear waves [3]. Inverse Q, also called specific attenuation or internal friction, is defined as $Q^{-1}=2\alpha(\omega)c(\omega)/\omega$, [2, Sect. 2.3], and if dispersion can be neglected, $c(\omega)=c_{0}$, it follows that $Q^{-1}\propto\omega^{y-1}$. Therefore, constant-Q behavior is the same as linearly increasing absorption.
Constant-Q behavior is especially apparent after correction for bias due to an additive constant in the expression for $\alpha(\omega)$, which may be due to geometrical spreading, (de-)focusing, or scattering, [3]. Similar results are reported for seismic reflection data from vertical wells [4] and constant-Q behavior up to about 100 Hz is also common in seismic survey data [5].
It is also recognized that $Q$ is nearly frequency-independent over one to two decades of frequency in many solids, [6]. This property was noted for metals and nonmetals, for both compressional and shear waves, and for frequencies in the Hz, kHz, and MHz range.
Unconsolidated sub-bottom sediments represent a very different medium, but the same linearly increasing absorption of compressional waves is observed above about 2 kHz [7]. Below that frequency the exponent is closer to two. Even in this case the value for the exponent, $y$, will depend on whether shear-wave mode conversion in the form of a constant term in the expression for $\alpha(\omega)$ is compensated for, [8].
@@
The medical ultrasound field has provided insight into the spatial scale where absorption takes place. Grinding liver tissue hardly changes the absorption, so it is apparent that the mechanisms operate ”on a level of organization smaller than that defined by cells, cell nuclei, and mitochondria”, i.e., size less than about 3 microns, [12]. Absorption in canned evaporated milk also follows a linear frequency law [13], indicating that casein micelles at the sub-micron size level play an important role. Similarly, absorption of compressional waves in blood in the 0.8-3 MHz range was due to the presence of proteins [14].
A related phenomenon is how cells or tissue react to shear excitation. The response often varies with $\omega^{\beta}$, where $\beta$ is in the range 0.1 to 0.5, [15], [16]. According to results cited in the Supplement, this corresponds to an absorption that varies with $y=\beta+1$ and therefore $y$ in this case will fall in the range 1.1 to 1.5.
The property that the exponent $y$ is near unity over a wide range of materials is nearly universal, both for compressional and shear waves, as noted in e.g. [17]. The exponent may be larger than unity, but never above 2, the viscous case, and rarely below 1. The temperature range of interest is approximately 0 to 40 C.
This paper is concerned with viscoelastic absorption, but there are other mechanisms which will not be discussed such as friction in cracks in rocks, fluid flow in micropores in consolidated sediments or bone, squirt flow in unconsolidated sediments, and nonlinearity.
An early study of absorption in erythrocytes suggested that “chemical or structural relaxation processes are probably responsible for the attenuation” [18]. It has also been proposed that in polymer-like materials, thermal energy causes continuous change in the “interactions between macromolecules” [17]. In cell rheology, the theory of soft glassy materials has been shown to be applicable, [15].
We also notice that chemical and structural relaxation in e.g., seawater are thermally activated processes. Thermal activation over potential barriers of structural defects also plays an important role in e.g. glasses and rocks. This gives the motivation to transform the multiple relaxation formulations from the usual frequency and time domains into the energy domain. This will provide new insights for the case of elastic wave propagation in acoustics and seismics. Surprisingly, a simple uniform distribution of activation energies corresponds to the universal linearly increasing absorption.
@@ II Background: Multiple relaxation
A relaxation process is characterized by a relaxation frequency, $\Omega$, or a relaxation time $\tau=1/\Omega$, and an absorption given by: ${\Omega\,\omega^{2}}/({\omega^{2}+\Omega^{2}})$. Attenuation increases with $\omega^{2}$ well below the relaxation frequency and is constant well above it. This expression can for instance be found from structural relaxation, [19], and from chemical relaxation, [20].
@@
$\alpha(\omega)=A_{0}\;\omega^{2}\int_{0}^{\infty}\frac{g_{\Omega}(\Omega)\,\Omega}{\omega^{2}+\Omega^{2}}\,\mathrm{d}\Omega,$ (2)
where $A_{0}$ is a constant. The weighting $g_{\Omega}(\Omega)$ in the relaxation integral has the form of a probability distribution function. The particular distribution given by $g_{\Omega}(\Omega)=\Omega^{y-2}$ will result in the power-law absorption of (1), [21, Sect. 3.241.2], [22]. This formulation does, however, not provide much insight to motivate why $g_{\Omega}(\Omega)$ should follow this particular power-law relation.
The integral of (2) can be transformed to be over relaxation times, $\tau$, by letting $g_{\Omega}(\Omega)=g_{\tau}(\tau)\;|\mathrm{d}\tau/\mathrm{d}\Omega|$:
@@
where the particular distribution $g_{\tau}(\tau)=\tau^{-y}$ will lead to the desired power-law absorption, [23]. This formulation may be easier to interpret as relaxation times may be related to length scales, $l=c\tau$, by means of the speed of propagation, $c$. Cell biomechanics may offer some insights, as power-law behavior of the shear modulus over about five decades of frequency is well documented. However, there seems not to be enough cell components in a hierarchy from the largest to the smallest to account for the observed power-law behavior, [24, 15, 25]. There is therefore limited physical insight to gain even from the formulation of (3).
@@ III Thermally activated relaxation
A closer look at the common mechanisms for intrinsic absorption is therefore warranted. In acoustics, they are viscosity, molecular thermal relaxation, heat conduction in monatomic gases, structural relaxation, and chemical relaxation, [26, Chap. 8], [2, Sect. 4.1]. Only the two last ones are relevant to properties of fluids and solids, our main interest in this paper. We therefore start by reviewing the well-established model for seawater, as an example of a medium where structural (also called segmental) relaxation and chemical relaxation dominate.
@@ III.1 The Arrhenius law
Absorption in seawater has three main components. Two of them are due to B(OH)3 and MgSO4 which both contribute two-state chemical equilibrium reactions. Their relaxation frequencies are in the kHz and tens of kHz range respectively. Relaxation time in these reactions is related to activation energy, $E_{a}$, according to the Arrhenius relation, [20]:
@@
where $T$ is absolute temperature, $k_{B}$ is Boltzmann’s constant, and $\tau_{0}$ characterizes the smallest time scale probed. It should be noted that in underwater acoustics, chemical relaxation is usually parameterized with temperature in Celcius rather than Kelvin, [27], obscuring the fact that these processes are thermally activated.
The most important contribution to absorption in water is from structural relaxation of H2O molecules. In distilled water a broken-down structure of clusters is dynamically changed by an incoming sound wave and relaxation takes place as clusters of different sizes interact. Based on a two-state energy model for H2O molecules, the Arrhenius law describes the transition rates, [19]. The relaxation frequency is in the THz range.
The Arrhenius law point to an activation energy landscape interpretation of chemical and structural relaxation taking place at the molecular or molecular cluster level, i.e. at the nanometer scale. For now, that only covers some of the cases mentioned in the Introduction, but let us anyway pursue the consequences of this view point.
@@ III.2 Transformation of relaxation integral
The validity of the Arrhenius relation is a key insight that allows us to transform the previous relaxation integrals. Equation (3) can be transformed by using (4) in combination with $g_{\tau}(\tau)=g_{E}(E_{a})/k_{B}T\cdot\;|\mathrm{d}E_{a}/\mathrm{d}\tau|$:
$\alpha(\omega)=\frac{A_{0}}{k_{B}T}\,\omega^{2}\int_{0}^{\infty}\frac{g_{E}(E_{a})\tau(E_{a})}{1+\omega^{2}\tau^{2}(E_{a})}\mathrm{d}E_{a},$ (5)
where $g_{E}(E_{a})$ is an energy distribution function. It should be noted that activation energy and its distribution function are physical parameters that lend themselves to interpretation much more than distributions of relaxation frequencies or relaxation times.
@@ III.3 Glass
Equation (5) is common for describing the effect of structural defects in glassy media, [28] with relaxation time given by the Arrhenius relation, [29]. A glass can be considered to be frozen in an energy basin with many local minima. The height of the barriers between them, and the energy difference between the minima will determine properties, [30]. Structural relaxation takes place due to perturbations of the landscape and this is the main cause of absorption. We are mainly concerned with normal temperatures (about 270 - 310 K), and then tunneling [31] can be neglected and the classical model with the thermal activation rate of (4) describes the relaxation.
@@ III.4 Energy distribution
$g_{E}(E_{a})={\Omega}\;g_{\Omega}(\Omega)={\tau}\;g_{\tau}(\tau).$ (6)
In the case of the power-law absorption of interest here, we have
$g_{E}(E_{a})={\Omega^{y-1}}.$ (7)
The most interesting case is the energy distribution corresponding to linearly increasing absorption, $y=1$. Surprisingly, that corresponds to a flat activation energy distribution.
A general distribution can be further investigated by combining (7) with (4), giving:
$g_{E}(E_{a})={\Omega_{0}^{y-1}}\mathrm{e}^{\frac{-E_{a}}{k_{B}T}(y-1)}.$ (8)
This result is plotted in Fig. 1. The most apparent feature is the flat distribution for $y=1$, as noted. In addition the case for $y>1$ is also interesting. It corresponds to an exponential energy distribution where lower energies are given more weight than higher energies.
Figure 1: Normalized energy distribution function of (8) plotted for power-law exponents $y$ in the range from 0.9 to 1.3 and for $\tau_{0}^{-1}=2\pi\cdot 10^{9}$.
So far we have given arguments for why (8) is valid for media where relaxation takes place at the nanometer scale. In the next section the size scale will be expanded.
IV Energy landscapes and biological materials
Interestingly, the energy landscape interpretation is used over a much larger range of scales in [33, Chap. 1]. Wales’s book starts by outlining three different fields: “The structure and dynamics of atomic and molecular clusters, the folding of proteins, and the complicated phenomenology of glasses are all manifestations of the underlying potential energy surface”

### v3

@@ Abstract
Absorption of elastic waves in complex media is commonly found to increase linearly with frequency, for both longitudinal and shear waves. This ubiquitous property is observed in media such as rocks, unconsolidated sediments, and human tissue. Absorption is due to relaxation processes at the level of atomic scales and up to the sub-micron scale of biological materials. The effect of these processes is usually expressed as an integral over relaxation frequencies or relaxation times. Here we argue that these processes are thermally activated. Unusual for ultrasonics and seismics, we can therefore transform the expression for absorption from the frequency or time domains to an integral over an activation energy landscape weighted by an energy distribution. The universal power-law property surprisingly corresponds to a flat activation energy landscape. This is the solution which maximizes entropy or randomness. Therefore the linearly increasing absorption corresponds to the energy landscape with the fewest possible constraints.
@@ I Introduction
The amplitude of elastic waves, i.e. both compressional and shear waves, undergoes attenuation which often follows a power-law in frequency, $\omega$:
@@
where ${u}$ is particle velocity and $x$ is distance. Attenuation, $\alpha$, has unit m-1 or Nepers/m and is also called inverse mean free path. The two mechanisms for attenuation are viscoelastic absorption leading to heating, and scattering of energy from inhomogeneities. This paper is only concerned with mechanisms for absorption, although medium properties may also be inferred from the scattering [1]. In complex media absorption often increases linearly with frequency, i.e. $y=1$ [2, Sect. 5.1], as the many examples that follow demonstrate.
In seismology, many nuclear-explosion and earthquake data sets in the range $10^{-3}$ to 10 Hz appear to have a constant quality factor, $Q$, for shear waves [3]. The unit-less inverse Q or attenuation per wavelength, (also called specific attenuation or internal friction), is [2, Sect. 2.3]:
$Q^{-1}=\frac{\alpha(\omega)\lambda(\omega)}{\pi}=2\frac{\alpha(\omega)c(\omega)}{\omega},$ (2)
where $\lambda$ is wavelength. When there is power-law absorption and dispersion is small, $c(\omega)\approx c_{0}$ [4], the inverse Q will be proportional to $\omega^{y-1}$. Therefore, linearly increasing absorption results in constant-Q behavior.
Constant-Q behavior is especially apparent after correction for bias due to an additive constant in the expression for $\alpha(\omega)$, which may be due to geometrical spreading, (de-)focusing, or scattering [3]. Constant-Q behavior is also reported for seismic reflection data from vertical wells [5] and such behavior is also common in seismic survey data up to about 100 Hz [6].
It is also recognized that $Q$ is nearly frequency-independent over one to two decades of frequency in many solids [7]. This property was noted for metals and nonmetals, for both compressional and shear waves, and for frequencies in the Hz, kHz, and MHz range.
Unconsolidated sub-bottom sediments represent a very different medium, but the same linearly increasing absorption of compressional waves is observed above about 2 kHz [8]. Below that frequency the exponent is closer to two. Even in this case the value for the exponent, $y$, will depend on whether shear-wave mode conversion in the form of a constant term in the expression for $\alpha(\omega)$ is compensated for [9].
@@
The medical ultrasound field has provided insight into the spatial scale where absorption takes place. Grinding liver tissue hardly changes the absorption, so it is apparent that the mechanisms operate ”on a level of organization smaller than that defined by cells, cell nuclei, and mitochondria”, i.e., size less than about 3 microns [13]. Absorption in canned evaporated milk also follows a linear frequency law [14], indicating that casein micelles at the sub-micron size level play an important role. Similarly, absorption of compressional waves in blood in the 0.8-3 MHz range was due to the presence of proteins [15].
The property that the exponent $y$ is near unity over a wide range of materials is nearly universal, both for compressional and shear waves, as noted in e.g. [16]. The exponent may in some cases be larger than unity, but never above 2, the viscous case, and rarely below 1. The temperature range of interest is approximately 0 to 40 C.
This paper is concerned with viscoelastic absorption, but there are other mechanisms which will not be discussed such as friction in cracks in rocks, fluid flow in micropores in consolidated sediments or bone [17], squirt flow in unconsolidated sediments [18], and nonlinearity.
An early study of absorption in erythrocytes suggested that “chemical or structural relaxation processes are probably responsible for the attenuation” [19]. It has also been proposed that in polymer-like materials, thermal energy causes continuous change in the “interactions between macromolecules” [16].
Here we claim that it can be justified to transform the multiple relaxation formulation into the energy domain for many of the materials just listed. This is a transformation into an energy landscape which is common for describing atomic and molecular clusters, glasses, and proteins.
We follow [20, Chap. 1] and define an energy landscape as either a potential energy surface or a free energy surface. Of these, the potential energy surface is the more fundamental as it gives the potential energy as a function of atomic or molecular coordinates. In complex media, this would be a function of many variables, which is unrealistic to know in detail. One then will replace it by a thermally averaged free energy surface which is described in terms of a smaller number of effective degrees of freedom. The distinction will not be important for us in the following, as we are only assuming certain statistical properties of the energy landscape, while not relating them to the microscopic structure.
Surprisingly, a simple uniform distribution of activation energies, i.e. maximal randomness, corresponds to universal linearly increasing absorption.
The paper starts with multiple relaxation formulations in frequency and time of power-law absorption. Then we introduce the Arrhenius law and use it to transform the relaxation formulation into the energy domain with an accompanying energy distribution. The question of whether energy landscapes also describe biological materials in then discussed. The paper ends with a discussion of non-Arrhenius behavior, the effect of band-limited power law absorption, and a comparison with a similar result as ours found in a completely different way in the field of soft glassy materials.
@@ II Background: Multiple relaxation
A single relaxation process is characterized by a relaxation frequency, $\Omega$, or a relaxation time $\tau=1/\Omega$, and an absorption given by:
$\alpha(\omega)=A\frac{\Omega\omega^{2}}{\omega^{2}+\Omega^{2}},$ (3)
where $A$ is a constant with unit inverse velocity. Attenuation increases with $\omega^{2}$ well below the relaxation frequency and is constant well above it. This expression can for instance be found from structural relaxation [21], and from chemical relaxation [22].
@@
$\alpha(\omega)=A_{0}\;\omega^{2}\int_{0}^{\infty}\frac{g_{\Omega}(\Omega)\,\Omega}{\omega^{2}+\Omega^{2}}\,\mathrm{d}\Omega.$ (4)
Following the terminology of [23], $A_{0}=2/c_{0}$, where $c_{0}$ is the equilibrium sound speed in the limit of zero frequency. The weighting $g_{\Omega}(\Omega)$ in the relaxation integral has the form of a probability distribution function. The particular distribution given by
$g_{\Omega}(\Omega)=K_{y}\;\Omega^{y-2},$ (5)
results in the power-law absorption of (1) [24, Sect. 3.241.2], [25], with a normalization factor $K_{y}$ with unit [sy-1]. This result was in fact already found in 1959 [26, Fig. 8]. This formulation does, however, not provide much insight to motivate why $g_{\Omega}(\Omega)$ should follow this particular power-law relation.
The integral of (4) can be transformed to be over relaxation times, $\tau=1/\Omega$, by letting $g_{\Omega}(\Omega)=g_{\tau}(\tau)\;|\mathrm{d}\tau/\mathrm{d}\Omega|$:
@@
where the particular distribution
$g_{\tau}(\tau)=K_{y}\tau^{-y},$ (7)
will lead to the desired power-law absorption [23]. This formulation may be easier to interpret as relaxation times may be related to length scales, $l=c\tau$, by means of the speed of propagation, $c$. A possible explanation for relaxation behavior can therefore be that there is a hierarchy of geometrical structures, from large $l$ to small $l$.
This interpretation has in particular been attempted in cell biomechanics which is a field where power-law behavior of the shear modulus over about five decades of frequency is well documented. Geometrical structures that may be invoked are the cell membrane, the actin cortex, the cytoskeleton etc, all at different length scales. It has been concluded however, that there are not enough cell components in a hierarchy from the largest to the smallest length scales to account for the observed power-law behavior, see [27, 28, 29] for the detailed argument. It seems therefore as if there is limited physical insight to gain even from the formulation of (6).
@@ III Thermally activated relaxation
A closer look at the common mechanisms for intrinsic absorption is therefore warranted. In acoustics, they are, according to [30, Chap. 8], [2, Sect. 4.1], viscosity, molecular thermal relaxation, heat conduction in monatomic gases, structural relaxation, and chemical relaxation. Viscous absorption and molecular thermal relaxation due to oxygen and nitrogen dominate in air, and the three first mechanisms therefore primarily characterize absorption in gases. Only the two last ones are relevant to properties of fluids and solids, our main interest in this paper. We therefore start by reviewing the well-established model for seawater, as an example of a medium where structural (also called segmental) relaxation and chemical relaxation dominate.
@@ III.1 The Arrhenius law
Absorption in seawater has three main components of the form of (3). Two of them are due to B(OH)3 and MgSO4 which both contribute two-state chemical equilibrium reactions. Their relaxation frequencies are in the kHz and tens of kHz range respectively. Relaxation time in these reactions is related to activation energy, $E_{a}$, according to the Arrhenius relation [22]:
@@
where $T$ is absolute temperature, $k_{B}$ is Boltzmann’s constant, and $\tau_{0}$ characterizes the smallest time scale probed. It should be noted that in underwater acoustics, chemical relaxation is usually parameterized with temperature in Celsius rather than Kelvin [31], obscuring the fact that these processes are thermally activated.
The most important contribution to absorption in water is from structural relaxation of H2O molecules. In distilled water a broken-down structure of clusters is dynamically changed by an incoming sound wave and relaxation takes place as clusters of different sizes interact. Based on a two-state energy model for H2O molecules, the Arrhenius law describes the transition rates [21]. The relaxation frequency is in the THz range.
The Arrhenius law points to an activation energy landscape interpretation of chemical and structural relaxation taking place at the molecular or molecular cluster level, i.e. at the nanometer scale. For now, that only covers some of the cases mentioned in the Introduction, but let us anyway pursue the consequences of this view point.
@@ III.2 Transformation of relaxation integral
The validity of the Arrhenius relation is a key insight that allows us to transform the previous relaxation integrals. Equation (6) can be transformed by using (8) in combination with $g_{\tau}(\tau)=g_{E}(E_{a})|\mathrm{d}E_{a}/\mathrm{d}\tau|$:
$\alpha(\omega)={A_{0}}\,\omega^{2}\int_{0}^{\infty}\frac{g_{E}(E_{a})\tau(E_{a})}{1+\omega^{2}\tau^{2}(E_{a})}\mathrm{d}E_{a},$ (9)
where $g_{E}(E_{a})$ is an energy distribution function. In this way, the Arrhenius law provides a link that allows us to transform the distribution of relaxation rates to a distribution of activation energies. These can then be interpreted in terms of the energy landscape.
@@ III.3 Glass
It was the use of (9) for describing the effect of structural defects in glassy media which inspired the work reported here. A glass can be considered to be frozen in an energy basin with many local minima. The height of the barriers between them, and the energy difference between the minima will determine properties [32]. Structural relaxation takes place due to perturbations of the landscape and this is the main cause of absorption. We are mainly concerned with typical terrestrial surface temperatures (about 270 - 310 K), and then tunneling [33] can be neglected and the classical model with the thermal activation rate of (8) describes the relaxation [34], [35], [36]
@@
In the glass field, the main interest is absorption as a function of temperature, and secondarily as a function of frequency. The model of (9) may either be used to find a distribution of activation energies, $g_{E}(E)$ that fits experimental data for absorption, or aspects of the distribution may be found from material properties. Common distributions are a gaussian [38], a gaussian weighted by a power law [39], and an exponential [40]:
$g_{E}(E_{a})=\frac{1}{E_{0}}\mathrm{e}^{\frac{-E_{a}}{E_{0}}},$ (10)
where $E_{0}$ is an energy related to the glass temperature.
The property that absorption follows a power-law as in (1) is not common to see in the description of glassy media, even though (9) is used extensively. One example that has power-law characteristics is found in [34], where they found an inverse Q that increases linearly with frequency, i.e. $y\approx 2$ in (2). This was found to correspond to a distribution of energies that falls off for higher energies, in general agreement with our result in the next section. Rather than look for power law absorption, it is more common to characterize absorption in glassy media by e.g. its peaks [40].
@@ III.4 Energy distribution
$g_{E}(E_{a})=\frac{\Omega}{k_{B}T}\;g_{\Omega}(\Omega)=\frac{\tau}{k_{B}T}\;g_{\tau}(\tau),$ (11)
where the relationship between activation energy and relaxation time is given by (8). In the case of the power-law absorption of interest here, given by (5) and (7), we have
$g_{E}(E_{a})=\frac{K_{y}}{k_{B}T}{\Omega^{y-1}}=\frac{K_{y}{\tau_{0}^{1-y}}}{k_{B}T}\;\mathrm{e}^{-\frac{E_{a}}{k_{B}T}(y-1)},$ (12)
where $\tau_{0}$ comes from (8). This is an exponential distribution, but due to the presence of the $y$, this result is more specific than (10).
The equation is plotted in Fig. 1 with the power-law exponent, $y$, as a parameter. The range of activation energies, $E_{a}/k_{B}$, is from 691 to 2072 K. Their ratio is 3 and so it corresponds to 3 decades of frequency according to (8). The lower and upper energy limits correspond to 100 MHz and 100 kHz respectively for a value of $\tau_{0}^{-1}=2\pi\cdot 10^{9}$.
The most interesting case is the energy distribution corresponding to linearly increasing absorption, $y=1$. Surprisingly, that corresponds to a flat activation energy distribution. This is not a feature which has received much attention in the energy landscape interpretation of e.g. glass. As $y>1$, the distribution falls off for higher energies and its mean value, assuming that the distribution is defined for all energies from 0 to infinity, is $k_{B}T/(y-1)$.
The energy distribution can be interpreted as a probability density function, $p(x)$, with proper normalization. The concept of Shannon entropy will aid in understanding its properties:
$H(p)=-\int_{a}^{b}p(x)\ln p(x)\mathrm{d}x.$ (13)
The density which maximizes entropy is a uniform or flat probability density function, when $a$ and $b$ are both finite. This interpretation of the flat energy landscape therefore corresponds to one with maximal randomness or one with the fewest possible constraints. Further, if the entropy is maximized over an infinite interval, $a=0$, $b=\infty$, with the constraint of a given mean value, the result is an exponential probability density function [41, Table 1]. Under that condition, the exponential solution in (11) also exhibits maximum randomness.
Figure 1: Normalized energy distribution function of (12) plotted for power-law exponents $y$ in the range from 0.9 to 1.3. The ratio of the upper and lower energy limits is 3 and therefore this plot corresponds to 3 decades of frequency variation.
So far we have given arguments for why (12) is valid for power-law media where relaxation takes place at the nanometer scale. In the next section the size scale will be expanded.
IV Biological materials and energy landscapes
Interestingly, the energy landscape interpretation is used over a much larger range of scales in [20, Chap. 1]. Wales’s book starts by outlining three different fields: “The structure and dynamics of atomic and molecular clusters, the folding of proteins, and the complicated phenomenology of glasses are all manifestations of the underlying potential energy surface”